OpenStudy (thesecret20111):

Solve these over the domain 0≤a≤2pi: 3cos^2a-5cosa-1=3sin^2a and 1/cos^2a+1-3tan^2a=0 I'm not sure what to do for either of these questions, can someone please help?

OpenStudy (dan815):

umm i think graph it

OpenStudy (dan815):

between that interval

OpenStudy (dan815):

oh nvm

OpenStudy (tkhunny):

Don't graph anything. Hint: \(\sin^{2}(x) = 1 - \cos^{2}(x)\)

OpenStudy (dan815):

do it like quadratic

OpenStudy (dan815):

they just want the solution to be in that interval

OpenStudy (thesecret20111):

Ok so like 3cos^2a-5cosa-1=3(1-cos^2(a))?

OpenStudy (dan815):

ya and then just put it on 1 side and finish like ud solve a quadratice equation

OpenStudy (thesecret20111):

Like 3cos^2a-5cosa-1-3-3cos^2(a)=0?

OpenStudy (dan815):

ya but +3cos^2

OpenStudy (dan815):

3cos^2a-5cosa-1-3+3cos^2(a)=0? 6cos^2a-5cosa-4=0

OpenStudy (dan815):